# Inflationary Krylov complexity

### v2

@@ Abstract
In this work, we have systematically investigated the Krylov complexity of curvature perturbation for the modified dispersion relation in inflation, using the algorithm in closed system and open system. Since many quantum gravitational frameworks could lead to this kind of modified dispersion relation, our analysis could be applied to the string cosmology, loop gravity, $\it e.t.c$. Following the Lanczos algorithm, we find the very early universe is an infinite, many-body, and maximal chaotic system. Our numerics shows that the Lanczos coefficient and Lyapunov index of the standard dispersion relation are mainly determined by the scale factor. As for the modified case, it is nearly determined by the momentum. In a method of the closed system, we discover that the Krylov complexity will show irregular oscillation before the horizon exits. The modified case will present faster growth after the horizon exists. As for the approach of an open system, we construct the exact wave function which is very robust only requiring the Lanczos coefficient proportional to $n$ (main quantum number). Based on it, we find the Krylov complexity and Krylov entropy could nicely recover in the case of a closed system under the weak dissipative approximation, in which our analysis shows that the evolution of Krylov complexity will not be the same with the original situation. We also find the inflationary period is a strong dissipative system. Meanwhile, our numerics clearly shows the Krylov complexity will grow during the whole inflationary period. But for the small scales, there will be a peak after the horizon exits. Our analysis reveals that the dramatic change in background (inflation) will significantly impact the evolution of Krylov complexity. Since the curvature perturbation will transit from the quantum level to the classical level. We could expect that the decoherence will highly impact the Krylov complexity during inflation.
@@
The holographic principle needs the conformal symmetry for the quantum field theory. Meanwhile, with the development of complexity in high-energy physics, it is natural to investigate the complexity in quantum systems. Ref. Aaronson:2016vto defined complexity as the minimal operations for a certain task, namely also dubbed as the computational complexity. Afterward, there were two main methods for calculating the complexity: (a) The geometric method of Nielsen et al Nielsen2006AGA ; Nielsen2006QuantumCA ; Dowling2008TheGO . (b) The method of ”Fubini-Study” distance on the related quantum states Chapman:2017rqy . To be more precise, these two main methods, especially the Nielson’s geometric method, are aimed to investigate the so-called circuit complexity, what they investigate is the minimal number of quantum operations from a reference state to a target state. The third method was called the Krylov complexity, Ref. Parker:2018yvk introduced the Krylov complexity as the measurement of operator growth with respect to Krylov basis, in which they proposed a universal operator growth in the chaotic system. Compared with the previous two methods, the definition of Krylov complexity is free of ambiguities since its definition is unique. For the other two methods, the complexity depends on the choice of quantum gates (the geometric method) or the metric (the method of ”Fubini-Study” distance). Once the quantum state is defined, one can construct the Krylov space in light of the Lanczos algorithm viswanath1994recursio , and one can calculate the Krylov complexity, Lanczos coefficient, and its corresponding entropy. Ref. Aguilar-Gutierrez:2023nyk has shown that one cannot make the Krylov and Nielsen notions of complexity compatible with each other, whereas under certain assumptions, one can allow for Krylov and ”Fuibini-Study” length to be proportional to each other Caputa:2021sib .
@@
Ref. Muck:2022xfc systematically developed the method for calculating the Krylov complexity in various orthogonal polynomials dubbed as the inner product. The Krylov complexity can be applied for the SKY model Parker:2018yvk ; Rabinovici:2020ryf ; Jian:2020qpp . One can also define a thermalized Krylov complexity Kar:2021nbm . Even the chaos can be understood as the delocalization in Krylov space Dymarsky:2019elm . By contrast, Ref. Rabinovici:2021qqt proposed the Krylov complexity is suppressed by the finite size which is in the integrable models. For further tests and numerical calculations, it can be implemented into the generalized coherent states Caputa:2021sib ; Patramanis:2021lkx , the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , CFT Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye , integrable models with saddle-point dominated scrambling Bhattacharjee:2022vlt , cosmological complexity Adhikari:2022oxr $\it e.t.c$. The very recent developments can be found in Refs. Erdmenger:2023wjg ; Patramanis:2023cwz ; Fan:2023ohh ; Hashimoto:2023swv ; Vasli:2023syq ; Domingo:2023kjr ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf . Especially, Ref. Camargo:2023eev involves examining the Krylov operator complexity within the simple quantum billiard model, which can be considered as the extension to finite temperature of Hashimoto:2023swv . Ref. Huh:2023jxt also has verified all the universal features of spread complexity within saddle dominated scrambling (”integrable” systems).
@@
One of the main motivations is that we aim to explore the Krylov complexity in the very early universe since the energy of the universe is not conserved due to the symmetry broken along the temporal part. It is naturally considering our universe as an open system. There are some attempts for using the complexity in cosmology, see Refs Adhikari:2021ked ; Choudhury:2020hil . To be more precise, one could consider that the universe will exchange information with the environment (outer of the universe). Following Refs. Bhattacharya:2023zqt ; Bhattacharjee:2022lzy , we will write the Hamilton of single field inflationary model with modified dispersion relation into one part that governs the closed and unitary system, another part will response the information exchanging with the environment. This kind of treatment can be considered as the Markovian environment in terms of Lindbladian evolution Lindblad:1975ef ; Gorini:1975nb . It should be noticed that the whole Hamilton is non-Hermitian under this Lindbladian evolution. From Ref. Socorro:2022aoz , it indicates that the gravity system (k-essence) is non-Hermitian. Therefore, it is quite natural to use the approach of open system.
@@
We organize the paper as follows. In Sec. II, we will give a brief introduction to Krylov complexity and the Lanczos algorithm, which is the method for dealing with closed system. In Sec. III, some fundamental knowledge about the modified dispersion relation and inflation will be given to readers. Sec. IV will give the calculation of Krylov complexity and its corresponding Krylov entropy and discuss some properties in inflation. In Sec. V, we will generalize the Lanczos algorithm into the open system and discuss the general properties of Krylov complexity in a weak dissipative system. Then, we will construct an exact wave function for calculating the Krylov complexity and Krylov entropy. Finally, the conclusion and outlook will be given in Sec. VI.
@@ II Krylov complexity and Lanczos algorithm
In this section, we will briefly introduce the concept and Krylov complexity and its corresponding Lanczos algorithm. It is convenient for working in the Heisenberg picture, first, we could have
@@
where $\mathcal{O}(t)$ is the operator depending on time in Heisenberg picture and $H$ is the Hamilton. Its corresponding solution could be written by
@@
where $k$ is the dimension of Krylov basis $\mathcal{O}_{n}$. In addtion, Ref. Parker:2018yvk has proposed that the Krylov complexity of the many-body system in the thermodynamic limit, the Lanczos coefficient $b_{n}$ will be conjectured that $b_{n}$ will linearly be approximated grow like,
@@ III.1 Modified dispersion relation
where $H=\frac{\dot{a}}{a}$, $\mathcal{R}=\psi+\frac{{H}}{\phi_{0}}\delta\phi$ , $z=\sqrt{2\epsilon}a$, $\epsilon=-\frac{\dot{H}}{H}=1-\frac{{H}^{\prime}}{H}$. Action (15) can be written in terms of the Mukhanov variable $v=z\mathcal{R}$,
@@ IV Krylov complexity of modified dispersion relation in closed system’s approach
In this section, we will compute the Krylov complexity in a closed system, which the method is following by Adhikari:2022oxr . The quantum state is a two-mode squeezed state as we previously discussed in Liu:2021nzx ; Li:2021kfq .
@@ IV.1 The evolution of $r_{k}$
$\begin{split}&\frac{dr_{k}}{d\eta}=\frac{k}{2}(f^{2}-1)\sin(2\phi_{k})+\frac{{z}^{\prime}}{z}\cos(2\phi_{k})\\ &\frac{d\phi_{k}}{d\eta}=\frac{k}{2}(f^{2}+1)+\frac{k}{2}(f^{2}-1)\cos(2\phi_{k})\coth(2r_{k})+\frac{{z}^{\prime}}{z}\sin(2\phi_{k})\coth(2r_{k})\end{split}$ (28)
@@
$\begin{split}\frac{10^{y}H_{0}}{\ln 10}\frac{dr_{k}}{dy}=\frac{k}{2}\bigg{[}(\frac{k_{ph}}{M})^{2\alpha}-1\bigg{]}\sin(2\phi_{k})+aH_{0}\cos(2\phi_{k})\end{split}$ (29)
$\begin{split}\frac{10^{y}H_{0}}{\ln 10}\frac{d\phi_{k}}{dy}=\frac{k}{2}\bigg{[}(\frac{k_{ph}}{M})^{2\alpha}+1\bigg{]}+aH_{0}\sin(2\phi_{k})\coth(2r_{k})+\frac{k}{2}\bigg{[}(\frac{k_{ph}}{M})^{2\alpha}-1\bigg{]}\cos(2\phi_{k})\coth(2r_{k})\end{split}$ (30)
@@ IV.2 Krylov complexity, Lanczos coefficient and Lyapunov index
Acting this Hamilton on the recursion relation, one can accordingly obtain that
@@
where it clearly indicates that the inflationary universe is an ideal chaotic system. One can derive the corresponding complexity based on (10),
@@
where we have used $\sum_{m=0}^{\infty}mz^{m}=\frac{z}{(1-z)^{2}}$. Then, (37). It should be pointed out that the two-mode squeezed state has the maximum complexity growth bound with the $SL(2,R)$ symmetric group structure Hornedal:2022pkc . In light of (12) and $\alpha=\frac{b_{n}}{n}$, it straightforwardly has
@@
Before investigating the Krylov complexity for a closed system, we simply analyze the features of Lyapunov index (38) and Lanczos coefficient (36). Explicit observations indicate that it will recover the standard dispersion relation as $f=1$. Through a simple algebra, one can obtain $\lambda=\frac{2}{\eta}$, $b_{n}=\frac{n}{\eta}$ and $K=\frac{1}{4k^{2}\eta^{2}}$ ($\eta$ is the conformal time) which is also consistent with Adhikari:2022oxr . At the very beginning of the universe with $\eta$ approaches $\infty$- ∞, one can explicitly obtain $K=0$. Especially, one could see that Krylov complexity is very tiny as $k\eta\ll 1$. For the late universe, the complexity will grow exponentially. These features are perfectly presenting the chaotic features of a dynamical system. However, once taking the impact of modified dispersion relation into account, it clearly shows that the Lyapunov index (38) and Lanczos coefficient (36) will depend on the momentum $k$ and $f$. We already know that its growth is linearly proportional to $n$. But, we still need to understand how $k$ and $a$ impact $b_{n}$ and $\lambda$.
@@
In Fig. 2, we found $b_{n}=\lambda$ as $n=2$ which is enough to manifest their changing trend. Keeping in mind that $\alpha$ will be larger than zero as $k_{\rm ph}>M$, and $k_{\rm ph}=k/a$. If we set $M=1$ (some certain energy scale), it means that $k$ should be larger than $a$ in Planck units. Meanwhile, $a=10^{y}$, if we set the maximal value of $y=3$, it requires that $k$ must be larger than this value. Through this simple analysis, that is the reason we set the range of $k$ from $1.1\times 10^{3}$ to $10^{6}$. The upper panel of Fig. 2 clearly indicates that $b_{n}=\lambda$ and $\lambda$ are nearly determined by $k$, which corresponds to the standard case. By contrast, the modified case shows that $b_{n}=\lambda$ and $\lambda$ are mainly determined by $a$. Thus, we can conclude that the geometrical effects of background will significantly impact the $b_{n}=\lambda$ and $\lambda$ for various quantum gravitational frameworks. The careful reader may notice that Fig. 2 does not include the case of $\alpha=3$ which we have confirmed that its behavior is almost the same with $\alpha=2$. Further, it gives us a new hint where the geometrical effects of background are essential for Krylov complexity.
@@
Figure 3: The plot of Krylov complexity (37), $y=\log_{10}(a)$ varies from -4 to 2. The ”horizon exit” is located at $y=-2.35$ (still in the very early universe).
Fig. 3 clearly indicates the varying trend of Krylov complexity with various values of $\alpha$ in a closed system. One can see that the complexity will grow faster compared with the standard case. To be more precise, the framework of quantum gravity, including the string cosmology, loop gravity $\it e.t.c$, will be more chaotic compared with the standard case. The obvious trend is that the complexity will grow faster as enhancing the value of $\alpha$. For completeness, we also plot the Krylov complexity before the horizon exits, and we find that the Krylov complexity will show the damping oscillation as $\alpha>2$ that is similar to our previous work Li:2023ekd . For the standard dispersion relation, one could see that the Krylov complexity is almost zero. Here, we make a conclusion that the complexity of quantum gravity including the Krylov complexity and circuit complexity will lead to the various patterns in the very early universe, whose evolution of circuit complexity can be found in Fig. 4 of Li:2023ekd . Generically, this damping oscillation is the criterion for assessing the quantum gravitational models and standard single-field inflationary models. The last thing is that the order of Krylov complexity is much larger than the circuit complexity under similar conditions.
There are other two interesting issues that need to be noticed which have been discussed in Ref. Adhikari:2022oxr . One thing is that the Krylov complexity equals the average number of particles in each mode $\left\langle\hat{n}_{k}\right\rangle=\left\langle\hat{n}_{-k}\right\rangle=\sinh^{2}r_{k}=K$. The average number density is proportional to the volume, thus one can explicitly obtain that $K\propto~{}volume$. In some sense, it demonstrates the CV conjecture in light of the holographic principle. Another issue is that the corresponding K-entropy can be defined by Barbon:2019wsy ,
@@
Let us make a summary of this section, we already use the Lanczos algorithm to calculate the Krylov complexity, Lanczos coefficient, Lyapunov index, and K-entropy. The Hamilton (23) what we implemented is the Hermitian and the wave function of the two-mode squeezed state is unitary. Consequently, the corresponding system is a closed system. However, the energy is not conserved for the universe since the temporal symmetry will be broken as the time-reversed. Thus, it is quite natural to extend the analysis into the open system.

### v4

@@ Abstract
In this work, we have systematically investigated the Krylov complexity of curvature perturbation for the modified dispersion relation in inflation, using the algorithm in closed system and open system. Since many quantum gravitational frameworks could lead to this kind of modified dispersion relation, our analysis could be applied to the string cosmology, loop gravity, $\it e.t.c$. Following the Lanczos algorithm, we find the very early universe is an infinite, many-body, and maximal chaotic system. Our numerics shows that the Lanczos coefficient and Lyapunov index of the standard dispersion relation are mainly determined by the scale factor. As for the modified case, it is nearly determined by the momentum. In a method of the closed system, we discover that the Krylov complexity will show irregular oscillation before the horizon exits. The modified case will present faster growth after the horizon exists. Since the whole universe is an open system, the approach of an open system is more realistic and reliable. Then, we construct the exact wave function which is very robust only requiring the Lanczos coefficient proportional to $n$ (main quantum number). Based on it, we find the Krylov complexity and Krylov entropy could nicely recover in the case of a closed system under the weak dissipative approximation, in which our analysis shows that the evolution of Krylov complexity will not be the same with the original situation. We also find the inflationary period is a strong dissipative system. Meanwhile, our numerics clearly shows the Krylov complexity will grow during the whole inflationary period. But for the small scales, there will be a peak after the horizon exits. Our analysis reveals that the dramatic change in background (inflation) will significantly impact the evolution of Krylov complexity. Since the curvature perturbation will transit from the quantum level to the classical level. We could expect that the decoherence will highly impact the Krylov complexity during inflation.
@@
The holographic principle needs the conformal symmetry for the quantum field theory. Meanwhile, with the development of complexity in high-energy physics, it is natural to investigate the complexity in quantum systems. Ref. Aaronson:2016vto defined complexity as the minimal operations for a certain task, namely also dubbed as the computational complexity. Afterward, there were two main methods for calculating the complexity: (a) The geometric method of Nielsen et al Nielsen2006AGA ; Nielsen2006QuantumCA ; Dowling2008TheGO . (b) The method of ”Fubini-Study” distance on the related quantum states Chapman:2017rqy . To be more precise, these two main methods, especially the Nielson’s geometric method, are aimed to investigate the so-called circuit complexity, what they investigate is the minimal number of quantum operations from a reference state to a target state. The third method was called the Krylov complexity, Ref. Parker:2018yvk introduced the Krylov complexity as the measurement of operator growth with respect to Krylov basis, in which they proposed a universal operator growth in the chaotic system. Compared with the previous two methods, the definition of Krylov complexity is free of ambiguities since its definition is unique. For the other two methods, the complexity depends on the choice of quantum gates (the geometric method) or the metric (the method of ”Fubini-Study” distance). Once the quantum state is defined, one can construct the Krylov space in light of the Lanczos algorithm viswanath1994recursio , and one can calculate the Krylov complexity, Lanczos coefficient, and its corresponding entropy. Ref. Aguilar-Gutierrez:2023nyk has shown that one cannot make the Krylov and Nielsen notions of complexity compatible with each other, whereas under certain assumptions, one can allow for Krylov complexity and ”Fuibini-Study” length to be proportional to each other Caputa:2021sib .
@@
Ref. Muck:2022xfc systematically developed the method for calculating the Krylov complexity in various orthogonal polynomials dubbed as the inner product. The Krylov complexity can be applied for the SKY model Parker:2018yvk ; Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk . One can also define a thermalized Krylov complexity Kar:2021nbm . Even the chaos can be understood as the delocalization in Krylov space Dymarsky:2019elm . By contrast, Ref. Rabinovici:2021qqt proposed the Krylov complexity is suppressed by the finite size which is in the integrable models. For further tests and numerical calculations, it can be implemented into the generalized coherent states Caputa:2021sib ; Patramanis:2021lkx , the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , CFT Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye , integrable models with saddle-point dominated scrambling Bhattacharjee:2022vlt , cosmological complexity Adhikari:2022oxr $\it e.t.c$. The very recent developments can be found in Refs. Erdmenger:2023wjg ; Patramanis:2023cwz ; Fan:2023ohh ; Hashimoto:2023swv ; Vasli:2023syq ; Domingo:2023kjr ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf . Especially, Ref. Camargo:2023eev involves examining the Krylov operator complexity within the simple quantum billiard model, which can be considered as the extension to finite temperature of Hashimoto:2023swv . Ref. Huh:2023jxt also has verified all the universal features of spread complexity within saddle dominated scrambling (”integrable” systems).
@@
One of the main motivations is to explore the Krylov complexity in the very early universe since the energy of the universe is not conserved due to the symmetry broken along the temporal part. It is naturally considering our universe as an open system. There are some attempts for using the complexity in cosmology, see Refs Adhikari:2021ked ; Choudhury:2020hil . To be more precise, one could consider that the universe will exchange information with the environment (outer of the universe). Following Refs. Bhattacharya:2023zqt ; Bhattacharjee:2022lzy , we will write the Hamilton of single field inflationary model with modified dispersion relation into one part that governs the closed and unitary system, another part will response the information exchanging with the environment. This kind of treatment can be considered as the Markovian environment in terms of Lindbladian evolution Lindblad:1975ef ; Gorini:1975nb . It should be noticed that the whole Hamilton is non-Hermitian under this Lindbladian evolution. From Ref. Socorro:2022aoz , it indicates that the gravity system (k-essence) is non-Hermitian. Therefore, it is quite natural to use the approach of open system.
@@
We organize the paper as follows. In Sec. II, we will give a brief introduction to Krylov complexity and the Lanczos algorithm, which is the method for dealing with closed system. In Sec. III, some fundamental knowledge about the modified dispersion relation and inflation will be given to readers. Sec. IV will give the calculation of Krylov complexity, its corresponding Krylov entropy and discuss some properties in inflation. In Sec. V, we will generalize the Lanczos algorithm into the open system and discuss the general properties of Krylov complexity in a weak dissipative system. Then, we will construct an exact wave function for calculating the Krylov complexity and Krylov entropy. Finally, the conclusion and outlook will be given in Sec. VI.
@@ II Krylov complexity and Lanczos algorithm
In this section, we will briefly introduce the concept of Krylov complexity and its corresponding Lanczos algorithm. It is convenient for working in the Heisenberg picture, first, we could have
@@
where $\mathcal{O}(t)$ is the operator depending on time and $H$ is the Hamilton. Its corresponding solution could be written by
@@
where $k$ is the dimension of Krylov basis $\mathcal{O}_{n}$. In addtion, Ref. Parker:2018yvk has proposed that the Krylov complexity of the many-body system in the thermodynamic limit, the Lanczos coefficient $b_{n}$ will be conjectured as follows,
@@ III.1 Modified dispersion relation
where $H=\frac{\dot{a}}{a}$, $\mathcal{R}=\psi+\frac{{H}}{\phi_{0}}\delta\phi$ , $z=\sqrt{2\epsilon}a$, $\epsilon=-\frac{\dot{H}}{H^{2}}=1-\frac{{\mathcal{H}}^{\prime}}{\mathcal{H}^{2}}$. Action (15) can be written in terms of the Mukhanov variable $v=z\mathcal{R}$,
@@ IV Krylov complexity of modified dispersion relation in closed system’s approach
In this section, we will compute the Krylov complexity in a method of closed system followed by Adhikari:2022oxr . The quantum state is a two-mode squeezed state as we previously discussed in Liu:2021nzx ; Li:2021kfq .
@@ IV.1 The evolution of $r_{k}$
$\begin{split}&-\frac{dr_{k}}{d\eta}=\frac{k}{2}(f^{2}-1)\sin(2\phi_{k})+\frac{{z}^{\prime}}{z}\cos(2\phi_{k})\\ &\frac{d\phi_{k}}{d\eta}=\frac{k}{2}(f^{2}+1)-\frac{k}{2}(f^{2}-1)\cos(2\phi_{k})\coth(2r_{k})+\frac{{z}^{\prime}}{z}\sin(2\phi_{k})\coth(2r_{k})\end{split}$ (28)
@@
$\begin{split}-\frac{10^{y}H_{0}}{\ln 10}\frac{dr_{k}}{dy}=\frac{k}{2}\bigg{[}(\frac{k_{ph}}{M})^{2\alpha}-1\bigg{]}\sin(2\phi_{k})+aH_{0}\cos(2\phi_{k})\end{split}$ (29)
$\begin{split}\frac{10^{y}H_{0}}{\ln 10}\frac{d\phi_{k}}{dy}=\frac{k}{2}\bigg{[}(\frac{k_{ph}}{M})^{2\alpha}+1\bigg{]}+aH_{0}\sin(2\phi_{k})\coth(2r_{k})-\frac{k}{2}\bigg{[}(\frac{k_{ph}}{M})^{2\alpha}-1\bigg{]}\cos(2\phi_{k})\coth(2r_{k})\end{split}$ (30)
@@ IV.2 Krylov complexity, Lanczos coefficient and Lyapunov index
where $\mathcal{H}$ denotes the part of closed system for Hamilton. Acting this Hamilton on the recursion relation, one can accordingly obtain that
@@
where it clearly indicates that the inflationary universe is an ideal chaotic system. It should be noted that the the Lanczos coefficient can also be determined by the interaction parameter Bhattacharyya:2023dhp . One can derive the corresponding complexity based on (10),
@@
where we have used $\sum_{m=0}^{\infty}mz^{m}=\frac{z}{(1-z)^{2}}$. It should be pointed out that the two-mode squeezed state has the maximum complexity growth bound with the $SL(2,R)$ symmetric group structure Hornedal:2022pkc . In light of (12) and $\alpha=\frac{b_{n}}{n}$, it straightforwardly has
@@
Before investigating the Krylov complexity within the method of closed system, we simply analyze the features of Lyapunov index (38) and Lanczos coefficient (36). Explicit observations indicate that it will recover the standard dispersion relation as $f=1$. Through a simple algebra, one can obtain $\lambda=\frac{2}{\eta}$, $b_{n}=\frac{n}{\eta}$ and $K=\frac{1}{4k^{2}\eta^{2}}$ ($\eta$ is the conformal time) which is also consistent with Adhikari:2022oxr . At the very beginning of the universe with $\eta$ approaches $\infty$- ∞, one can explicitly obtain $K=0$. Especially, one could see that Krylov complexity is very tiny as $k\eta\ll 1$. For the late universe, the complexity will grow exponentially. These features are perfectly presenting the chaotic features of a dynamical system. However, once taking the impact of modified dispersion relation into account, it clearly shows that the Lyapunov index (38) and Lanczos coefficient (36) will depend on the momentum $k$ and $f$. We already know that its growth is linearly proportional to $n$. But, we still need to understand how $k$ and $a$ impact $b_{n}$ and $\lambda$.
@@
In Fig. 2, we found $b_{n}=\lambda$ as $n=2$ which is enough to manifest their changing trend. Keeping in mind that $\alpha$ will be larger than zero as $k_{\rm ph}>M$, and $k_{\rm ph}=k/a$. If we set $M=1$ (some certain energy scale), it means that $k$ should be larger than $a$ in Planck units. Meanwhile, $a=10^{y}$, if we set the maximal value of $y=3$, it requires that $k$ must be larger than this value. Through this simple analysis, that is the reason we set the range of $k$ from $1.1\times 10^{3}$ to $10^{6}$. The upper panel of Fig. 2 clearly indicates that $b_{n}$ and $\lambda$ are nearly determined by $k$, which corresponds to the standard case. By contrast, the modified case shows that $b_{n}$ and $\lambda$ are mainly determined by $a$. The careful reader may notice that Fig. 2 does not include the case of $\alpha=3$ which we have confirmed that its behavior is almost the same with $\alpha=2$. Further, it gives us a new hint where the geometrical effects of background are essential for Krylov complexity.
@@
Figure 3: The plot of Krylov complexity (37), $y=\log_{10}(a)$ varies from -4 to 2. The ”horizon exit” is located at $y=-2$ (still in the very early universe). We have set $M=H_{0}=1$.
Fig. 3 clearly indicates the varying trend of Krylov complexity with various values of $\alpha$ in a method of closed system. One can see that the complexity will grow faster compared with the standard case. To be more precise, the framework of quantum gravity, including the string cosmology, loop gravity $\it e.t.c$, will be more chaotic compared with the standard case. The obvious trend is that the complexity will grow faster as enhancing the value of $\alpha$. For completeness, we also plot the Krylov complexity before the horizon exits, and we find that the Krylov complexity will show the damping oscillation as $\alpha>2$ that is similar to our previous work Li:2023ekd . For the standard dispersion relation, one could see that the Krylov complexity is almost zero. Here, we make a conclusion that the complexity of quantum gravity including the Krylov complexity and circuit complexity will lead to the various patterns in the very early universe, whose evolution of circuit complexity can be found in Fig. 4 of Li:2023ekd . Generically, this damping oscillation is the criterion for assessing the quantum gravitational models and standard single-field inflationary models. The last thing is that the order of Krylov complexity is much larger than the circuit complexity under similar conditions.
There are other two interesting issues that need to be noticed which have been discussed in Ref. Adhikari:2022oxr . One thing is that the Krylov complexity equals the average number of particles in each mode $\left\langle\hat{n}_{k}\right\rangle=\left\langle\hat{n}_{-k}\right\rangle=\sinh^{2}r_{k}=K$. The average number density is proportional to the volume, thus one can explicitly obtain that $K\propto~{}volume$. In some sense, it illustrates the CV conjecture in light of the holographic principle. Another issue is that the corresponding K-entropy can be defined by Barbon:2019wsy ,
@@
Let us make a summary of this section, we already use the Lanczos algorithm to calculate the Krylov complexity, Lanczos coefficient, Lyapunov index, and K-entropy. The Hamilton (23) what we implemented is the Hermitian and the wave function of the two-mode squeezed state is unitary. The method what we implement is for the closed system. However, the energy is not conserved for the universe since the temporal symmetry will be broken as the time-reversed. Thus, it is quite natural to extend the analysis into the open system.